# Efficient Indirect LLM Jailbreak via Multimodal-LLM Jailbreak

### v2

@@ 4.3.1 Image-text Semantic Matching
Figure 3: Cross-class generalization for LLM-jailbreak. For example, the first row illustrates that we generate JBtxt on the “Violence" and then apply it to jailbreak other harmful classes.
4.3.2 Cross-class Generalization
In real-world jailbreak scenarios, it is desirable for the JBtxt generated for a specific harmful class to generalize and effectively jailbreak other harmful classes.
To evaluate this cross-class generalization capability, we conduct an experiment where JBtxt is generated based on a specific harmful class and subsequently tested on all remaining classes. For instance, after generating a JBtxt using Class 1’s training set $D_{train}^{1}$, we evaluate its ASR on the testing sets of the other classes: $D_{test}^{2}$, $D_{test}^{3}$, …, $D_{test}^{9}$. This setup allows us to assess whether a JBtxt optimized for one class can successfully jailbreak across different harm classes.
The results of the cross-class generalization experiment are presented in Fig.3. We observe that certain classes, such as “Financial Crimes", “Property Crimes", and “Cyber Crimes" exhibit strong generalizability. Specifically, the JBtxt generated from any harmful class can effectively jailbreak these three classes. In contrast, classes like “Hate" and “Suicide and Self-Harm" prove challenging to generalize. Notably, even when the JBtxt is generated from these two classes themselves, the testing ASR remains low, as reflected by the diagonal entries in Fig.3.
Additionally, we observe that cross-class generalization exhibits distinct correlations between different classes. Certain pairs of classes demonstrate notably stronger correlations than others. For example, the JBtxt generated for the class “Unlawful Violence" can effectively jailbreak the class “Fake Info," while the JBtxt derived from “Weapons Crimes" shows strong efficacy in jailbreaking “Drug Crimes." This phenomenon implies that to improve the ASR for a given class, it may be advantageous to leverage harmful queries not only from that specific class but also from its correlated classes. This approach is particularly valuable when collecting sufficient queries for a specific harmful class is difficult. In future work, we aim to investigate the underlying factors driving these correlations among harmful classes.
5 Conclusion

### v3

@@ Discussion
Cross-class Generalization
In real-world jailbreak scenarios, it is desirable for the JBtxt generated for a specific harmful class to generalize and effectively jailbreak other harmful classes. We conduct an experiment to evaluate the cross-class generalization capability of our approach. We find that certain classes, such as “Financial Crimes”, “Property Crimes”, and “Cyber Crimes” exhibit strong generalizability. In contrast, classes like “Hate” and “Suicide and Self-Harm” prove challenging to generalize. The detailed results are provided in the supplementary material.
Conclusion
@@
Acknowledgments
This work was supported in part by Natural Science Basic Research Program of Shaanxi (Program No.2025JC-JCQN-075).